OpenStudy (anonymous):

During his first year of life,Avery gained 12 3/4 lb. During his second year of life, Avery gained 40% less weight than he did during his first year of life. How much weight did Avery gain during his second year of life? Express your answer as a mixed number in simplest form. lb

Vocaloid (vocaloid):

you know, if you don't understand something, you can always ask... anyway, convert 12 3/4 to a decimal... what is 3 divided by 4?

OpenStudy (anonymous):

ohh ok thxs

Vocaloid (vocaloid):

answer my question please

Vocaloid (vocaloid):

what is 3 divided by 4?

OpenStudy (anonymous):

0.75

Vocaloid (vocaloid):

good, so 12 3/4 as a decimal is 12.75 now multiply that by 0.6

OpenStudy (anonymous):

7.65

Vocaloid (vocaloid):

good, now convert that back into a mixed number

OpenStudy (anonymous):

how?

Vocaloid (vocaloid):

look at the decimal part .65 = 65/100 now reduce that fraction

OpenStudy (anonymous):

opps

OpenStudy (anonymous):

iim so confused

Vocaloid (vocaloid):

start with 65/100 now divide the numerator and the denominator by 5

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

65/100=0.65/5=0.13

Vocaloid (vocaloid):

nope what is 65/5?

OpenStudy (anonymous):

13

Vocaloid (vocaloid):

good so 65/100 = 13/20 so 7.65 is equal to 7 13/20 as a mixed number and that's our final answer

OpenStudy (anonymous):

ok thxs